Ticket Number: HEA-NOC/20080513-1 Ticket Status: UPDATE
Ticket Type: unscheduled Resolver: Dublin City Council
Ticket Opened: 20080513 09:22 UTC+1 Problem Start: 20080512 21:05 UTC+1
Ticket Update: 20080523 12:37 UTC+1 Problem End:
Site/line: DCU, DIT, Digital Depot, RIA, SVUH, UCD
Dublin City Council fibre has been damaged due to a fibre cut near St.
Stephens green. Multiple clients effected
Loss of resiliency for SVUH (St. Vincents), Belfield, DIT Aungier , Bolton
St, DCU Mater Dei and Eccles Street
Total outage for Royal Irish Academy the Oireachtas and RIFTFS
20080513 09:11 COC Contacted DCC who are aware of the problem and are
working to fix it. They are trying to have the Fibre repaired by the end
20080513 17:07 BN DCC have yet to effect repairs on the circuit, but are
continuing to work to restore service. It is hoped that connectivity will
be restored by tomorrow morning at 09:00, but this is not guaranteed.
20080514 09:23 BN DCC have attempted a temporary work-around for this
outage, but no sites had connectivity restored. A full repair process is
underway, but due to the location of the fibre break the estimated
restoration date is Friday 16th May. This ticket will be updated with any
changes to this timeline and details communicated to the relevant parties.
20080515 10:04 BN DCC did a lot of work last night, so the revised
estimate of restoration of service is between 17:00 - 18:00 this evening,
Thursday 15th. This will be closely monitored by HEAnet and all affected
clients kept informed.
20080516 11:15 BN Overnight work by DCC has repaired fibres, but only the
Oireachtas has returned to service. Testing will be carried out today to
attempt to restore the remaining clients.
20080516 11:59 BN Oireachtas and RIA circuits have had connectivity
restored. RIFTFS is still down and troubleshooting is continuing.
20080516 18:12 BN RIFTFS came back up at 17:59. Connectivity has been
confirmed. Bolton St, Mater Dei and Eccles St have not had resiliency
restored, DCC are continuing to work on this.
20080519 09:48 TG Spoke with DCC, one of our 3 pairs to TCD is still not
up, they are planning a further outage on either Tuedday or Wednesday
night to get this issue fully resolved
20080523 12:31 BN DCC wish to carry out some further remedial work on
Monday 26th May to finally re-close the ring disrupted by the initial
fibre break. This may cause outages to RIA, RIFTS and the House of the
Oireachtas as fibre pairs are moved and/or respliced.
This ticket can be monitored at http://www.hea.net/tickets/20080513-1
HEAnet Limited. Registered in Ireland, No. 275301.